A radio station broadcasts in the Western Cape on 99.5 fm. The radio waves transmitted from this station have a frequency of 99,5MHz.
Calculate the wavelength of these waves.
1
Expert's answer
2020-04-20T10:42:27-0400

The wavelength of a 99.5  MHz radio signal is


"\lambda=\frac{c}{f}=\frac{3\cdot 10^8}{99.5\cdot 10^6}=3.015\ m"